MONT 104N – Modeling the Environment Paper Assignment 2 – Priorities for the Next Four Years due: November 16, 2012

In the second debates day on Wednesday, November 14, we will hear a collection of opinions on what should be President Obama’s main priority in his second term. These opinions come from a New York Times Room For Debate blog published on November 7:

          http://www.nytimes.com/roomfordebate/2012/11/07
        /a-to-do-list-for-president-obamas-next-four-years

This should be all entered as one web address; it’s broken over two lines to fit on the page here. There is also a link to this on the course homepage in connection with the second debate day.

This paper assignment has two main goals:

1. Imagine you personally got to enter a debate like the one presented in the New York Times. Try to articulate what you think should be the main priority for the President and the country over the next four years and how you would advise President Obama to try to address it. This may or may not have a strong environmental component, but if your goal is primarily non-environmental, you should try to say why you think action on some of the environmental challenges we face is less important than your primary goal (or why addressing environmental challenges can be postponed until other problems have been addressed).

2. Everyone should work to write this paper along the lines of the templates from They Say, I Say. The first section should be a fair summary of two or three of the positions presented in the New York Times blog. You can choose which of them you want to include, and the choice should depend on what you want to say. For instance, based on what you think should be the main priority for the next four years, you might include one position with which you agree strongly and one with which you disagree strongly, then frame your response in agreement and opposition to those. Or, if you believe that none of the positions taken in the discussion reflect what you would see as the correct major priority, you might frame your response as a rejection of all of them. Be sure you think about how all of the “moves” presented in parts 2 and 3 of They Say, I Say might be useful here.

Important Note: It is perfectly OK with me if you want to share a rough draft of your essay with other students in the class, or with me, to get comments and suggestions. “Testing out” your ideas that way is often an important part of getting them down on paper in an effective form.

Prepare your essay using Microsoft Word and submit as a .doc file attachment to this email address:

                           jlittle@holycross.edu

Your assignment should arrive no later than 5:00pm on Friday, November 16. Aim for about 5 double-spaced pages, but slightly over or under will not be a problem.

Include a complete listing of all sources you consulted in a numbered References section at the end, including the individual entries from the Times blog that you are drawing on for your argument. If you take a direct quotation from one of the Times blog entries, or from another source, identify it with an in-line reference of the form “(number, page)” where the number refers to the list of references at the end. Include a page number for printed sources; if you use online sources, that part can be omitted since there is usually not anything corresponding to a physical page number anyway(!)